OpenStudy (anonymous):

5x + 3 = 2x + 15 solve for x

OpenStudy (apoorvk):

subtract '2x' from both sides first, what do you get?

OpenStudy (apoorvk):

@tnicole22

OpenStudy (anonymous):

12

OpenStudy (jackellyn):

You want to get the x alone. Get all the x-es on one side and the numbers on the other side Then divide by the coefficient on x. Example: 2x-1=3x+3 2x-3x=3+1 -x = 4 x = 4 Try to do that with your problem.

mathslover (mathslover):

\[\huge{5x+3=2x+15}\] \[\huge{5x+3-2x=15}\] \[\huge{3x+3=15}\] \[\huge{3x=15-3}\] \[\huge{3x=12}\] \[\huge{x=\frac{12}{3}=4}\]